Run Coordination Meetings:Fall2018 Run

From GlueXWiki
Jump to: navigation, search

Fall 2018 GlueX Run

Meeting General Agenda

Connect Information

To connect from the outside by phone please do the following

1.) To join via a Web Browser, go to the page [1]

2.) To join via Polycom room system go to the IP Address: ( and enter the meeting ID: 660743227.

3.) To join via phone, use one of the following numbers and the Conference ID: 660743227.

4.) The moderator code is 8394

Dial Either +1 408 740 7256 or +1 888 240 2560 for US or Canada

Specific instructions for connecting from JLab's Hall D conference room:

  • Turn polycon on if necessary (do it before turning the computer on)
  • With the polycon, place a call at
  • Press # to enable the polycom keypad, then enter the meeting id: 660743227 and #
  • You may have to unmute the microphone: #*4
  • Turn the computer on if needed

Beam and Hall configuration

  • Beam energy expected:
    • 11.640 GeV (same value as Fall 2016 to Spring 2018) until Nov. 15th.
    • Then 10.3 GeV until Dec. 10th.
    • Then 9.0 GeV until Dec. 19th.
  • Solenoid on at 1350A for the 3 energies. ComCal commissioning with solenoid off. Any other field changes should be avoided if possible.
  • CW Beam current:
    • Range: 1 nA-1.5 μA. 250 MHz frequency.
    • Chopper slit shared with Hall A (Hall A to start run on Sept. 12th)
    • Standard production running expected at 530 nA on 1.4·10-4 R.L. radiator (17 μm diamond).
    • RSAD limit for most of the run, excluding high intensity tests: 600 nA on 1.6·10-4 R.L. radiator.
  • FFB on or Position Slow Lock on (FFB doesn't operate below 50nA).
  • Radiators:
    • Goniometer radiators:
      • Diamond JD70-104 (17 μm - 1.4·10-4 R.L., 7×7 mm²) - wider rocking curve - Tested in Spring 2018: Yielded reasonable polarization.
      • Diamond JD70-105 (47 μm - 3.9·10-4 R.L., 7×7 mm²) - new crystal - good rocking curve
    • 40 μm Al radiator (44.9·10-5 R.L.)
    • V-wire (for TAC run).
    • Amorphous
      • 1.64 µm Al (1.86·10-5 R.L., see study)
      • 10 µm Al (11.2·10-5 R.L.)
      • 30 µm Al (33.7·10-5 R.L.)
  • Tagger quadrupole on (-4.2 A).
  • Collimator hole: Both 5 mm diameter and 3.4 mm diameter. Default running: 5.0 mm, including with 17 μm diamond.
  • Targets:
  • Miscellaneous:
    • DIRC will be installed
    • Transition Radiation Detector in the Hall
    • Muon chamber will not be in the hall (location interferes with DIRC)

General Information

This document describes the run plan for the Hall D Fall 2018 run.

  • Schedule for the run:
  1. Aug. 15th - 21st: Electron beam restoration
  2. Aug. 21st - Nov. 26th: GlueX-I run (E=11.64 GeV).
  3. Nov. 26th - Nov. 28th: Beam energy change/ComCal installation (Solenoid will be ramped down but kept cold)
  4. Nov. 29th - Dec. 09th: Primex tests. (E=10.3 GeV)
  5. Dec. 10th - Dec. 12th: Linac energy change.
  6. Dec. 13th - Dec. 19th: Primex tests. (E=9.0 GeV)
  1. Every Mondays: 12h of beam study. Every Tuesdays, if necessary: 8h of RF recovery.
  • Accelerator overall plan and priorities:
  1. Deliver physics beam to 4 halls. (Hall A to starts later, on Sept 12th)
  • Experiment: GlueX E12-06-102 Production
  • Planned tests (Accelerator, related to Hall D):
    • Further tests of nA BPM and stripline BPM ?
    • Test fast raster beam envelope measurement with larger x-rastering. Need detailed plan, including time estimate
  • Planned tests (Hall D):
    • At 11.6 GeV:
      • Align JD70-105 (47 μm - 3.9·10-4 R.L., 7×7 mm²) (Hovanes) and take enough data to achieve ~1% statistic on rho polarimetry (less than a shift at 190 nA).
      • Test fast raster beam envelope measurement with larger x-rastering. (R. Jones) Need detailed plan, including time estimate
      • High-intensity GlueX L1 and DAQ tests (A. Somov, S. Furletov)
      • ToF trigger test for Muon chamber. With solid target. (D. Lawrence/E. Smith) Need detailed plan, including time estimate
      • TRD test (parasitic) (S. Furletov, L. Penchev)
    • At 10.2 and 9 GeV:
      • DIRC commissioning (J. Stevens) Need detailed plan, including time estimate
      • ComCal commissioning (no B-field, no DC on) (A. Somov) [2] [3] [4] [5]
      • Strait track run (L. Penchev)
      • High-intensity GlueX L1 and DAQ tests (A. Somov, S. Furletov)
  • Expected Staffing and responsibilities:
    • The Run Coordinator oversees the commissioning:
      • Aug. 15th-Aug. 21st, 7 days: Alexandre Deur (accelerator restoration)
      • Aug. 22nd-Aug. 28th, 7 days: Alexandre Deur
      • Aug. 29th-Sept 4th, 7 days: Cristiano Fanelli
      • Sept. 5th-Sept. 11th, 7 days: Alexandre Deur (accelerator down for repair)
      • Sept. 12th-Sept. 18th, 7 days: Alexander Austregesilo
      • Sept. 19th-Sept. 25th, 7 days: Thomas Britton
      • Sept. 26th-Oct. 2nd, 7 days: Naomi Jarvis
      • Oct. 3rd-Oct. 9th, 7 days: Stuart Fegan
      • Oct. 10th-Oct. 16th, 7 days: Colin Gleason
      • Oct. 17th-Oct. 23rd, 7 days: Ben. Zihlmann (Deputy PDL: A. Deur)
      • Oct. 24th-Oct. 31st, 7 days: Daniel Lersch
      • Oct. 31th-Nov. 7th, 7 days: Naomi Jarvis
      • Nov. 7th-Nov. 14th, 7 days: Alexander Austregesilo
      • Nov. 14th-Nov. 20th, 7 days: Alexandre Deur
      • Nov. 21st-Nov. 27th, 7 days: Ben. Zihlmann
      • Nov. 28st-Dec. 4th, 7 days: Wenliang Li
      • Dec. 5th-Dec. 11th, 7 days: Liping Gan
      • Dec. 12th-Dec. 20th, 8 days: Jeorg Reinhold
    • The Physics Division Liaison (PDL) and Hall D Work Coordinator/Safety Warden (M. Stevens) verify that the proper safety rules are followed. PDLs are:
      • Benedikt Zihlmann (except during Oct. 17th-Oct. 23rd: A. Deur to be PDL).
    • The analysis coordinator (Alexander Austregesilo) oversees the off-line diagnostic of the different systems and analysis of the data.
    • Shifts: 2 persons per shift.
  • Coordination Activities:
    • The RC will attend the MCC meetings at 7:45am and 8:00am and the weekly MCC meeting (Wednesday, 1:30pm). The PDL will attend the weekly MCC meeting.
    • Daily run meetings broadcasted on Bluejeans will occur at 8:45am in the counting house conference room. The Run Coordinator will be responsible for organizing and chairing the meeting.
  • Useful links:


  1. Verify electron beam quality and establish photon beam (assume 1 day).
    • Ion Chambers trip threshold must be set using Al. radiators
    • Tune electron beam with radiator retracted
      • Radiator (amorphous and diamond) must be retracted. Collimator should be in blocking position.
        • Typical levels are (for radiator retracted, collimator fully blocking, electron beam current 50 nA)
          • RAD102_P1 ~ 10 mrad/hr (tagger area, gammas, between tagger and dump)
          • RAD102_P2 < 0.1 mrad/hr (tagger area, gammas, on the ground below the goniometer).
          • RAD102_P3 ~ 0.3 mrem/hr (tagger area, neutrons, near electronics racks)
          • RAD508_P1 < 0.1 mrad/hr (collimator cave, gammas).
          • RAD508_P2 ~ 0. mrad/hr (Hall. Neutrons).
    • Establish good photon beam
    • Make sure tagger quadrupole is on and at -4.2 A. (It can be turned on only once the Hall is n power/beam permit.)
      • Insert 2*10-5 RL radiator. Call MCC and ask to mask the FSD and turn off beam each time a radiator is moved
        • Typical levels are (for 2×10-5 radiator, electron beam current 50 nA)
          • RAD102_P1= 35 mrad/hr (tagger area, gammas, between tagger and dump)
          • RAD102_P2= < 0.5 mrad/hr (tagger area, gammas, on the ground below the goniometer)
          • RAD102_P3= ~2 mrem/hr (tagger area, neutrons, near electronics racks)
          • RAD508_P1= ~10 mrad/hr (collimator cave, gammas).
          • RAD508_P2 < ~ 0. mrad/hr (Hall. Neutrons).
    • Verify beam position and envelope.
      • Ask MCC to set beam position on 5C11b to x=-0.5 mm and y=+1.0 mm.
      • Ask MCC for 5C11 and 5C11B HARP. Then do our radiator HARP scan, see instruction. If it looks good, proceed to next step. Otherwise, inform MCC.
      • Insert profiler (no need for masking FSD when an object is inserted in the photon line), run with 10×10-4 radiator and 50 nA. Beam on the profiler should be at x=0.0±0.5cm; y=0.0±0.5cm. If not, inform MCC.
    • Radiation level study (1h):
      • Run 5 min without radiator (Call MCC and ask to mask the FSD each time the radiator is moved). Beam current: 50nA. Repeat at 100 nA and 200 nA
      • Run 5 min with the 10-4 radiator (Call MCC). Beam current: 50nA. Repeat at 100 nA and 200 nA
      • Run 3 min with the 3*10-4 radiator (Call MCC). Beam current: 50nA. Repeat at 100 nA and 200 nA
    • PS detector threshold scan. (A. Somov, 1h)
    • Microscope yield study (row-by-row scan). (Richard/Jim McIntyre, 1h)
      • Call Richard Jones 20 min before the start of this task. Call Alex Somov to set-up the PS trigger.
      • Insert 5mm collimator.
      • Procedure to follow
      • Any type of beam quality will do.
      • 10-4 radiator, 10-3 convertor, 80 nA beam current.
    • Photon flux optimization and Active Collimator calibration. (20 + 20 mins). Call Hovanes.
      • re-Insert the 10-4 radiator (Call MCC). Use about 100 nA beam current. This will provide the Act. Col. nominal positions for its high gain values.
      • Determine maximum photon beam transmission by doing 2D scans using the collimator x/y motions (Hovanes. 1h)
      • Repeat procedure with 2*10-5 radiator (Call MCC) and 50 nA beam current. This will provide the Act. Col. nominal positions for its low gain values.
      • Record on the white board what is the beam position on 5C11b and Active Collimator for optimal photon transmission, both for low and high Act. Col. gain.
      • Until Hovanes determines to optimal 5C11b and Active Collimator position, keep profiler in and beam locked (MCC's duty) on the nominal 5C11b and x=0,y=0 profiler position. Once Active Collimator nominal beam position is known, retract profiler and lock the beam (MCC) on the active collimator.
  2. Detector and beamline checkout.
    • Reestablish L1 trigger Alex. Somov. 2h).
    • Reestablish DAQ (Sergei, 2h).
    • General detector calibration (Sean Dobbs. 1h). Use 10-4 radiator.
      • Run 15 min (solid time) in Long Mode (Mode 10).
      • Long mode (Mode 10) w/ FCAL LED on 1-10Khz (Mark D.)
      • Run 15 min (solid time) in Production Mode (Mode 9).
      • CDC HV scan Procedure to follow (Naomi Jarvis, 4h)
  3. High beam current systematic test (8h. Can be done before or after diamond alignment, as most convenient for the alignment)
    • Insert 10 μm Al. radiator (1.12*10-4 RL, on amorphous rad. ladder). Take 200M events (standard prod.) at 680 nA beam current. Monitor carefully rates in wire chambers.
    • Insert 40 μm Al. radiator (4.49*10-4 RL, on goni.). Take 200M events (standard prod.) at 170 nA beam current. Monitor carefully rates in wire chambers.
  1. Align 17 μm diamond (JD70-104) for the 4 polarization configurations (Hovanes. 16h).
  2. Gluex data production
    • Physics production data with 17 μm diamond, 530 nA, 5 mm hole, 75 μm TPol convertor. Ratio of para./transverse/45o/135o data:25/25/25/25. 300M events per configuration. 2h runs and no more, as it may be hard to correct for calibration drifts. Switch polarization every 300M events. Time to switch: 10min. DAQ start run overhead: 3min.
    • Amorphous run. Every set of 4 diam. orientation runs. 150M events. 170 nA on 40 μm Al. radiator (4.49*10-4 RL, on goni.). Time to switch from pol. to unpol.: 10min. time to switch from unpol. to pol.: 30min.
    • Log in run information in the shift summary (make a run list). For each production run, log useful comments on DAQ comment window, see [6]
    • Tagger quadrupole on
    • Harp scans once a day: 15 min.
    • If commissioning and analysis are successful, once a day, a beam envelope measurement by rastering on Act. Col. ??min
    • Empty target run every 2 week (2h: 1- h to empty/fill back the target, 1h running). Standard production current.
    • Long mode 10 min run. Once a week (20 min).
  3. Align JD70-105 (47 μm - 3.9·10-4 R.L., 7×7 mm²) for 2 or 4? polarization configurations (Hovanes 12h) and take enough data to achieve ~1% statistic on rho polarimetry (8h at 190 nA). Decide if it is better to run on this diamond.
  4. Two TAC runs at the beginning and end of 11.6 GeV run. May require other halls to be off due to bleedthrough (Hovanes, Alex Somov 8/h run.)
  5. Luminosity scans (J. Stevens)?
  6. Some physics production data with thin diamond and 3.4 mm hole 11h (1 cycles of runs). More data may be taken if beam stability and assessed polarization warranty it.
  7. ToF runs for muon chamber? Need solid target. Need detailed plan, including time estimate
  8. Parasitic TRD run
  9. During 10.2 GeV and 9.0 runs:
    • Re-establish good photon beam. Ion Chamber thresholds to be set each time electron beam energy changes.
    • Set PS field to rescaled value?
    • Realign one of the diamond?
    • DIRC commissionning (J. Stevens) Need detailed plan, including time estimate
    • ComCal commissioning (A. Somov) Need detailed plan, including time estimate
    • Gather data to be compared to CLAS results. (J. Stevens) Intersperesed with above commissioning. Need detailed plan, including time estimate
    • L1 and DAQ high luminosity tests (A. Somov/S. Furletov)
    • Luminosity scans? (J. Stevens)
    • Straight track run (16h end of 10.2 GeV period) (Lubomir)/ComCal test (A. Somov?)
      • Ramp down solenoid magnet (8h)
      • Take straight track run data (Lubomir)
      • ComCal test (no B-field, no DC on)
    • ToF runs for muon chamber? Need solid target. Need detailed plan, including time estimate
    • Parasitic TRD runs

Hall D situation room (current run plan)

Hall D counting room white board
GoogleDoc Spreadsheet of Trigger Statistics

Runplan time charts

Fall 2018 time chart:

Fall18 week1.jpg

Fall 2018 commissioning/run time chart. Week 1.
The horizontal scale is the time during which beam is available (in other words, it would be the real time with 100% of beam efficiency).
The task time estimates given on the plot have 100% contingency.

Fall18 week2.jpg

Run time chart. Week 2.

Fall18 week3.jpg

Run time chart. Week 3.

Fall18 week4.jpg

Run time chart. Week 4.

Fall18 week5.jpg

Run time chart. Week 5.

Fall18 week6.jpg

Run time chart. Week 6.

Fall18 week7.jpg

Run time chart. Week 7.

Fall18 week8.jpg

Run time chart. Week 8.

Fall18 week9.jpg

Run time chart. Week 9.

Fall18 week10.jpg

Run time chart. Week 10.

Low energy run (10.3 and 9 GeV)

Fall18 week11.jpg

Run time chart. Week 11.

Fall18 week12.jpg

Run time chart. Week 12.

Fall18 week13.jpg

Run time chart. Week 13.

Accelerator status

Hot Checkout

  1. Hot Check OUT Status 100% ready ; 1% checked ; 0% not ready

Readiness for Fall 18 run

  • Solenoid (Eugene/Beni)
    • Ramped to 100A, then 800A, and then back to 0A.
    • Ready
    • Waiting for accelerator firm delivery prediction to start ramping it up.
  • Surveys (Tim/Mark)
      • Diamonds: done
      • TAC: done
      • LH2 target: done
      • (DIRC, Monday 08/20)
  • DAQ (Sergei).
    • Running and ready.
    • Good sample of cosmic data was taken
  • Trigger (Alex S.)
    • Need to test new software for FADC
  • Beam line
    • Diamond status (Tim/Hovanes).
    • Tagger hodoscope (A. Somov)
    • Tagger microscope (Beni/Richard Jones)
      • All electronic has been turned on. Richard is checking the signals.
    • Goniometer (Hovanes).
    • Collimator, Active collimator, profiler, luminosity monitor (Hovanes/Alexandre Deur).
      • Act. Col. baselines have been zeroed mid July. Checked on 08/15: they have not changed significantly.
      • luminosity monitor usual location is incompatible with ComCal location.
    • PS system (Alex Somov)
    • PS magnet (and sweeping magnet and tagger quads) will be ramped up by A. Deur when he cools the target.
    • Polarimeter (Sebastian Cole/M. Dugger).
      • Ready
    • Target (Tim/Alexandre D.)
      • LH2 Target was cooled down (test). Target ready and HCO done.
      • Now target is warmed again, to be cooled on Aug. 21st. Cooling the target takes 6h. It will remains empty (so that there is no need to man the counting room) and be filled when needed.
      • Solid target status
    • TAC (Hovanes/Alex Somov)
      • Ready.
      • TDC now working.
        • To be cured (oven) after the Fall 2018 run.
        • During TAC runs to be inserted on beam line with 2cm offset to avoid damaged area
    • Slow control (Hovanes).
      • Control for DIRC and ComCal to be done, when detectors are installed.
    • RF (Elton/Beni)
      • Sergey F. found a flaky cable in the interface box. It is working for now but Sergey will contact Fernando about fixing this.
    • Radiation monitors (Alexandre D.)
      • Ready
    • Halo Monitors (Hovanes)
      • Ready
  • Detector systems
    • ST (Mark Ito)
      • Ready
    • FDC (Lubomir)
      • One channel has failed and is turned off.
      • Ready
    • CDC (Beni)
      • Ready
    • BCal (Elton/Mark D.)
      • Ready
    • FCal (Collin Gleason)
      • Ready. HCO done.
    • TOF (Mark Ito/Beni/Alex. Ostrovidov)
      • Ready
    • DIRC (J. Stevens)
      • Water test successful.
    • ComCal (Alex Somov)
  • Safety documents (Beni):
    • RSAD and ESAD ready.
  • Phones
    • System experts need to start carrying the on-call (charged) phones
  • Hall Status/safety (Mark/Tim)
  • Other topics?

Run Planning Meetings Fall 2018 Running

Run Summaries